Help : Q Login for >e l Q Homel m. × E-Mail-sen x | e) Review A, × 1 Digital Rex 8298 Ch. 4, Elas 50% 01/21/19 ce ntity De One of the problems with the basic elasticity formula is that you get different elasticity values depending on whether you assume the price falls from $8 to $6, or whether you assume the price rises from $6 to $8. The question belo llustrates how the midpoint formula corrects for this problem. 8th attempt See Hint Part 1 (1 point) Feedback The price elasticity of demand if the price of a pint falls from $8 to $6 is VIEW SOLUTION K0219> 9 OF 19
Answer #1

Price Ed = % change in quantity demanded / % change in price

% change in price = (P2 - P1)/P1 x 100 = (6 - 8)/8 x 100 = - 25

% change in quantity demanded = (Q2 - Q1)/Q1 x 100 = (5 - 3)/3 x 100 = 66.67

Price Ed = 66.67/(-25) = - 2.67